# Depth-Structured Music Recurrence: Budgeted Recurrent Attention for Full-Piece Symbolic Music Modeling

### v1

@@ Abstract
Long-context modeling is essential for symbolic music generation, since motif repetition and developmental variation can span thousands of musical events. However, practical composition and performance workflows frequently rely on resource-limited devices (e.g., electronic instruments and portable computers), making heavy memory and attention computation difficult to deploy. We introduce Depth-Structured Music Recurrence (DSMR), a recurrent long-context Transformer for full-piece symbolic music modeling that extends context beyond fixed-length excerpts via segment-level recurrence with detached cross-segment states, featuring a layer-wise memory-horizon schedule that budgets recurrent KV states across depth. DSMR is trained in a single left-to-right pass over each complete composition, akin to how a musician experiences it from beginning to end, while carrying recurrent cross-segment states forward. Within this recurrent framework, we systematically study how depth-wise horizon allocations affect optimization, best-checkpoint perplexity, and efficiency. By allocating different history-window lengths across layers while keeping the total recurrent-state budget fixed, DSMR creates depth-dependent temporal receptive fields within a recurrent attention stack without reducing compute depth. Our main instantiation is a two-scale DSMR schedule that allocates long history windows to lower layers and a uniform short window to the remaining layers. Experiments on the piano performance dataset MAESTRO demonstrate that two-scale DSMR provides a practical quality–efficiency recipe for full-length long-context symbolic music modeling with recurrent attention under limited computational resources.
I Introduction
@@
Long-context symbolic music modeling is a canonical setting where long-range temporal structure is indispensable yet computationally demanding [21, 35]. Musical coherence depends on context spanning many bars or entire pieces (e.g., motif repetition, phrase development, and global form), while practical creation and on-device experimentation often operate under limited compute and memory budgets. This combination makes efficient long-context modeling central to realistic symbolic music generation.
Transformer self-attention is powerful but scales quadratically with sequence length [31], becoming a dominant bottleneck for long musical streams. Despite many efforts to improve Transformer efficiency, practical full-length, long-context training and inference for symbolic music on consumer-grade GPUs remains largely unsupported by existing models.
To make full-piece context feasible under limited computational resources, we propose Depth-Structured Music Recurrence (DSMR). DSMR adopts Transformer-XL style recurrence with stop-gradient KV memory as a practical backbone for long-context symbolic music modeling. Unlike Transformer-XL, which typically uses a fixed memory length and thus discards earlier segments, DSMR exposes each composition in a single left-to-right pass, carrying recurrent states forward within the same piece segment by segment, so that every token can condition on long-range causal history spanning the entire piece. This also differs from the widely used windowed training in language modeling, which truncates sequences into fixed-length excerpts and treats them as independent samples.
Once recurrence mitigates time-axis costs, the key question shifts from how much history is available to how history is exposed across depth. Standard recurrent Transformers implicitly treat cross-segment memory as a uniform, fixed-window per-layer cache. In contrast, recurrent KV memory can be viewed as defining each layer’s temporal receptive field–the amount of past context accessible to recurrent attention at that depth. This viewpoint leads to a expressive mechanism:
DSMR uses a layer-wise memory-horizon schedule for recurrent KV caching. Under a fixed recurrent-state budget, DSMR assigns different history-window lengths $m^{(\ell)}$ across depth (including $m^{(\ell)}=0$), thereby creating depth-dependent temporal receptive fields within a recurrent attention stack. This lets different layers specialize to different effective context ranges, enabling the model to represent both short- and long-range dependencies without increasing compute depth.
Importantly, DSMR’s memory mechanism is not compute-depth pruning: all $L$ layers are still computed within each segment. DSMR instead structures cross-segment information flow by shaping layer-wise receptive fields, raising a central question: under fixed resource budgets, does long-context performance depend on the identity of the layers that access history, or on the overall receptive-field allocation across depth?
We study DSMR in a budgeted setting where the total recurrent-state budget is fixed. Within this regime, our final model uses a two-scale DSMR schedule for recurrent KV caching: lower layers are assigned long history windows, while the remaining layers share a uniform short window. This structural prior induces depth-dependent temporal receptive fields under a fixed budget, while keeping training and implementation straightforward.
To understand the DSMR design space, we additionally study two analysis-only variants. First, we consider binary-horizon DSMR (selective retention), where only $B$ layers have nonzero horizons (others have 0), to test whether particular layer identities are crucial when a fixed depth budget is enforced. Second, we introduce memory usage gating as an interpretability probe: learned gates modulate the historical KV contribution during training, and we test whether gate magnitudes provide transferable guidance for selective retention.
Our contributions are threefold. First, we introduce Depth-Structured Music Recurrence (DSMR), a recurrent long-context framework for full-piece symbolic music modeling that exposes each composition in a single left-to-right pass and carries cross-segment states forward to preserve full-piece causal history, while enforcing layer-wise temporal receptive fields via per-layer horizons $\{m^{(\ell)}\}$ under a fixed recurrent-state budget. Second, we propose two-scale DSMR as our final model, allocating long horizons in lower layers and a uniform short horizon in the remaining layers, which is a strong default for long-context symbolic music modeling. Third, we provide a variant analysis, showing (i) strong structural substitutability under binary-horizon DSMR at fixed depth budgets, and (ii) that gate preference does not provide a reliable, transferable predictor of gains under intervention. Together, these contributions yield a practical quality–efficiency recipe for long-context symbolic music modeling with recurrent attention, enabling fast training with substantially reduced memory compared to a full-attention reference while retaining competitive perplexity.
@@
The remainder of this paper is organized as follows. Section II reviews related work on recurrent Transformers, long-context mechanisms, and symbolic music modeling. Section III presents DSMR and its two-scale instantiation. Section IV describes the experimental setup, including the matched-budget protocol and the model variants. Section V reports the main results, including additional model variants, followed by efficiency evaluation. Section VI discusses mechanisms, scope conditions, and implications for designing recurrent long-context music models. Appendix A reviews stateful recurrence and establishes notation.
II Related Work
Symbolic music is inherently long-context: realistic pieces often exceed typical context limits, and salient phenomena such as motif repetition, thematic development, and sectional form can span tens of thousands of events. Scaling Transformers to such long sequences is therefore particularly important for music models, yet challenging because vanilla self-attention incurs quadratic time and memory in sequence length. Prior work tackles the cost of Transformer attention through two complementary directions: subquadratic long-context mechanisms (e.g., sparse/linear/state-space alternatives) and streaming-friendly stateful recurrence that reuses representations across segments. We review these lines and their budgeted memory designs, with an emphasis on symbolic music modeling under resource constraints.
II-A From Quadratic Attention to Subquadratic Long-Context Mechanisms
A large literature reduces attention cost by introducing sparsity patterns or approximations over the time axis. Representative approaches include sparse/block attention [8], sliding-window and global-token schemes [3], and sparse graph-style attention [36]. Kernelized or linearized attention targets near-linear scaling by rewriting attention in a feature space [22, 9]. Beyond attention, state space sequence models (SSMs) replace attention with a fixed-size recurrent state updated in linear time; Mamba introduces input-dependent (selective) state updates and achieves strong long-sequence modeling efficiency [18]. More recently, neural long-term memory modules aim to combine fast recurrent-style updates with attention-level accuracy, e.g., Titans and the accompanying MIRAS blueprint [2, 1]. At the system level, recent LLM efforts also propose highly optimized sparse attention mechanisms for long-context training and inference, such as DeepSeek Sparse Attention (DSA) [14]. Overall, attention optimization in standard Transformers offers a rich design space for reducing long-context cost.
II-B Recurrent Transformers and Budgeted Stateful Memory
@@
Complementary to attention-efficiency methods (e.g., sparsity), long-context modeling can also be achieved via stateful recurrence, where hidden states from previous segments are cached and reused as additional attention context. Transformer-XL introduced segment-level recurrence with cached activations and relative positional encoding, enabling dependencies beyond a fixed window while reducing context fragmentation[12], and is particularly attractive for streaming due to segment-bounded compute.
@@
Building on Transformer-XL-style recurrence, prior work has explored budgeted mechanisms that extend history under limited memory. Compressive Transformer summarizes older states into a lower-resolution buffer to trade fidelity for longer coverage [29]. Other approaches introduces explicit persistent memories, e.g., Recurrent Memory Transformer (RMT) carries learned memory slots across segments [6], and ERNIE-Doc augments recurrence with extended history and a retrospective feed to strengthen cross-segment information flow [16]. These approaches exemplify recurrent-attention-specific efforts beyond vanilla KV caching.
II-C Layer-Wise Memory Optimization
@@
Recurrent Transformers carry key/value states across segments as a persistent KV cache. This cache is often redundant across depth, with substantial inter-layer similarity in cached states [24, 38]. Motivated by this, inference-time studies on non-recurrent Transformers examine layer-wise KV-cache utility, showing it is heterogeneous across depth and making uniform budgeting suboptimal: SqueezeAttention assigns layer-wise KV budgets based on sensitivity [33], LAVa derives dynamic layer (and head) budgets from information-loss considerations [30], and MiniCache exploits depth-wise redundancy by compressing highly similar KV states across adjacent middle-to-deep layers [24].
Depth heterogeneity also appears in recurrent stacks, where long-range access is provided by stateful memories carried across segments. Rae et al. [28] show that Transformer-XL-style recurrence can retain comparable language-modeling performance with substantially fewer long-range memories, suggesting that long-range capacity need not be uniformly allocated across layers.
@@
II-D Long-Context Symbolic Music Modeling under Resource Constraints
@@
A central goal in symbolic music generation is to capture long-range structure and full-piece coherence, since motifs, harmonic plans, and section-level returns can span thousands of events. To this end, music-specific Transformer variants introduce structure-aware mechanisms for efficiency and long-term coherence, including improved relative attention [21], fine–coarse attention over salient vs. nonsalient bars [35], hierarchical/segmented whole-piece factorization [34, 7], and long-context pretraining for symbolic-music foundation models [27]. Alternatives to token-level autoregression, such as diffusion and latent-unit approaches, are also emerging for extreme sequence lengths [26], and LLM-style large-scale symbolic models further raise expectations on long-range coherence and musicality [32].
Yet in practice, training pipelines still often depend on fixed-length segmentation or bounded-context truncation, or on hierarchical/cascaded generation with fixed-scope local modeling, rather than end-to-end full-song training. This parallels common LLM pretraining pipelines based on truncation/packing with EOS concatenation to improve compute and memory utilization [39]. Recent LLM work suggests such fragmentation can harm data integrity: Ding et al. [15] report that “concatenate-then-chunk” may unnecessarily break documents that would otherwise fit within the context window, degrading coherent-content modeling.
Our focus on full-length music training targets this methodological gap. By reducing structural breaks and distribution shifts induced by cropping boundaries, full-song training can shift learning beyond local texture and short-term continuity toward global form and multi-minute dependencies. This supports higher-level musical generation behaviors such as more stable section planning, more consistent motif repetition, and more coherent overall orchestration/arrangement. Moreover, by improving recurrent memory mechanisms, we aim to make long-context training feasible on consumer-grade GPUs, lowering the barrier to iteration and deployment in creator-centric workflows where composition and production are frequently performed on portable computers and constrained hardware.
III Method
@@
We propose Depth-Structured Music Recurrence (DSMR), a training-time design for stateful recurrent attention that enforces layer-wise temporal receptive fields by assigning depth-dependent memory horizons under a fixed overall recurrent-state budget. DSMR keeps the Transformer backbone unchanged and only reshapes where along depth long-range history is retained. Our final model is a two-scale DSMR: long horizons in lower layers and a uniform short horizon elsewhere.
III-A DSMR’s recurrent backbone
@@
We adopt a Transformer-XL style segment-level recurrence. Each layer caches past key/value states and reuses them as additional attention context for subsequent segments. Moreover, unlike Transformer-XL which discards older segments, although the specific layer(s) that carry long memory may differ across model variants, DSMR ensure that every model retains at least one full-length (maximum-horizon) memory layer so that long-range context is never deterministically discarded. For completeness, details of Transformer-XL recurrence and its approximation to full self-attention are provided in Appendix A. Fig. 1 illustrates the segment-level recurrence with cached history and discarded context beyond the horizon.
@@
III-B Training-time sampling and segmentation
@@
III-C Full-attention approximation of a vanilla Transformer as a reference model
Transformer-XL enforces a fixed uniform horizon $m$ by retaining only the most recent $m$ positions (Eq. (14)), so attention to history is limited to a sliding window [12]:
$(K_{<t},V_{<t})\approx\bigl(\mathrm{Last}_{m}(K_{1:(t-1)}),\ \mathrm{Last}_{m}(V_{1:(t-1)})\bigr).$ (1)
Appendix A-A provides the exact stateful recurrent attention formulation for Transformer-XL, including memory concatenation, truncation, and stop-gradient across segments.
To approximate full-length context, DSMR extend Transformer-XL by setting the horizon to the maximum available prefix length for all layers, yielding a full-memory recurrent model.
Because DSMR uses truncated backpropagation through time (BPTT) with a stop-gradient across segments, it induces a biased approximation to full self-attention over the entire prefix, since the cached states are produced by parameters from earlier optimization steps. Let $\theta_{t}$ denote the parameters at training step $t$. Then the recurrent attention at step $t$ uses history states computed by earlier parameters,
$(K_{<t},V_{<t})=\bigl(K_{<t}(\theta_{<t}),\ V_{<t}(\theta_{<t})\bigr),$ (2)

### v2

# arXiv:2602.19816v2[cs.SD] 02 Mar 2026

@@ Abstract
Long-context modeling is essential for symbolic music generation, since motif repetition and developmental variation can span thousands of musical events, yet practical workflows frequently rely on resource-limited hardware. We introduce Depth-Structured Music Recurrence (DSMR), a training-time design that learns from complete compositions end to end by streaming each piece left-to-right with stateful recurrent attention and distributing layer-wise memory horizons under a fixed recurrent-state budget. Our main instantiation, two-scale DSMR, assigns long history windows to lower layers and a uniform short window to the remaining layers. On the MAESTRO piano performance dataset, two-scale DSMR matches a full-memory recurrent reference in perplexity (5.96 vs. 5.98) while using approximately 59% less GPU memory and achieving roughly 36% higher throughput. Variant analyses further show strong layer substitutability under binary-horizon schedules: performance depends primarily on total allocated memory rather than which layers carry it.
keywords:
Symbolic music generation , Long-context modeling , Recurrent attention , Streaming training , Depth-Structured Music Recurrence
journal: Knowledge-Based Systems
1 Introduction
@@
Musical coherence depends on context that spans many bars or entire compositions: motif repetition, phrase development, and sectional form can stretch across thousands of symbolic events Huang et al. [2019], Yu et al. [2022]. At the same time, practical music creation and on-device experimentation often operate under tight compute and memory budgets. Efficient long-context modeling is therefore central to realistic symbolic music generation.
Transformer self-attention is effective for sequence modeling but scales quadratically with length Vaswani et al. [2017], making it a dominant bottleneck for long musical streams. Although subquadratic alternatives, e.g., sparse attention Child et al. [2019], Beltagy et al. [2020], linear attention Katharopoulos et al. [2020], and state-space models Gu and Dao [2024], reduce per-step cost, they do not address a separate, equally important problem: training-time context fragmentation. Standard language-model pipelines truncate long sequences into fixed-length excerpts and optimize them as independent samples Ding et al. [2024]. This windowed regime is convenient for batching, yet it systematically fragments piece-level structure in symbolic music, where salient phenomena routinely span an entire composition.
Segment-level recurrence, as introduced by Transformer-XL Dai et al. [2019], offers a natural remedy: by caching key/value (KV) states across segments, the model can condition on long-range history without recomputing the preceding segments. However, standard recurrent Transformers treat cross-segment memory as a uniform, fixed-window cache at every layer. This uniform allocation is wasteful: not all depths require the same amount of history, and retaining full-length memory at every layer is prohibitively expensive on consumer-grade hardware.
In this paper, we propose Depth-Structured Music Recurrence (DSMR), a training-time design that makes full-piece end-to-end learning feasible under constrained resources. DSMR streams each composition through stateful recurrent attention and distributes layer-wise memory horizons under a fixed total recurrent-state budget, so that different layers access different amounts of past context. Our main instantiation is a two-scale DSMR schedule: lower layers retain long history windows while the remaining layers share a uniform short window. On the MAESTRO piano performance dataset Hawthorne et al. [2019], two-scale DSMR matches a full-memory recurrent reference in perplexity (5.96 vs. 5.98) while using approximately 59% less GPU memory and achieving roughly 36% higher throughput. Variant analyses further reveal strong layer substitutability under binary-horizon (selective retention) schedules, indicating that performance depends primarily on how much total memory is allocated rather than which specific layers carry it, and that learned memory-usage gates, while exhibiting non-uniform depth preferences, do not reliably transfer as guidance for choosing memory-carrying layers.
Our contributions are threefold. First, we introduce DSMR as a training-time framework that enables end-to-end learning on complete symbolic music compositions by streaming each piece left-to-right with stateful recurrent attention, where the carried cross-segment states are budgeted through a layer-wise horizon schedule so that full-piece training remains feasible under limited compute. Second, we propose two-scale DSMR as the main instantiation: long horizons are allocated to lower layers and a uniform short horizon to the remaining layers, keeping recurrent-state cost within a fixed budget without reducing model depth. Third, we provide systematic variant analyses that characterize the DSMR design space, establishing that depth-wise memory redistribution outperforms binary layer selection at matched budgets, and that gate-derived layer rankings do not reliably transfer under structural intervention.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 presents the DSMR framework and its two-scale instantiation. Section 4 describes the experimental setup. Section 5 reports the main results and variant analyses. Section 6 discusses implications, scope conditions, and limitations.
2 Related Work
We review prior work along four axes: subquadratic attention mechanisms, stateful recurrence with budgeted memory, layer-wise memory optimization, and long-context symbolic music modeling.
2.1 Subquadratic Long-Context Mechanisms
A large body of work reduces attention cost through sparsity or approximation. Representative approaches include sparse/block attention Child et al. [2019], sliding-window and global-token schemes Beltagy et al. [2020], sparse graph-style attention Zaheer et al. [2020], and kernelized/linearized attention Katharopoulos et al. [2020], Choromanski et al. [2021]. Beyond attention modifications, selective state-space models such as Mamba achieve linear-time sequence modeling with input-dependent state updates Gu and Dao [2024], and hybrid architectures such as Titans augment attention with persistent memory modules Behrouz et al. [2025]. At the system level, sparse attention mechanisms optimized for long-context LLM efficiency have also been proposed DeepSeek-AI [2025]. These methods reduce per-step cost but are complementary to the training-time concern addressed by DSMR: even with efficient attention, standard pipelines still fragment long sequences into fixed-length excerpts, discarding piece-level structure.
2.2 Recurrent Transformers and Budgeted Stateful Memory
@@
Complementary to subquadratic attention, long-context modeling can be achieved via stateful recurrence, where hidden states from previous segments are cached and reused as additional attention context. Transformer-XL introduced segment-level recurrence with cached activations and relative positional encoding, enabling dependencies beyond a fixed window while reducing context fragmentation Dai et al. [2019]. Its segment-bounded compute makes it particularly attractive for streaming workloads, and it serves as the recurrent backbone adopted by DSMR. However, standard Transformer-XL uses a uniform, fixed-length memory window at every layer and deterministically discards states beyond that window; DSMR departs from this design by assigning layer-wise horizons so that different depths retain different amounts of history.
@@
Building on Transformer-XL-style recurrence, several budgeted mechanisms extend history under limited memory. Compressive Transformer summarizes older states into a lower-resolution buffer, trading fidelity for longer temporal coverage Rae et al. [2020]. Recurrent Memory Transformer (RMT) carries learned memory slots across segments Bulatov et al. [2022], and ERNIE-Doc augments recurrence with extended history and a retrospective feed to strengthen cross-segment information flow Ding et al. [2021]. These approaches budget memory through compression or fixed-size slots, whereas DSMR budgets memory by redistributing per-layer horizons under a fixed total.
An inherent concern with cached cross-segment states is staleness: as the number of recurrent steps grows, earlier cached states become increasingly outdated relative to current parameters. While staleness has been studied in pipelined and asynchronous optimization Narayanan et al. [2019], Zhang et al. [2016], its impact on training dynamics in full-piece end-to-end music streaming—where a single composition may span over 30 recurrent steps—has not been previously examined.
2.3 Layer-Wise Memory Optimization
@@
Recurrent Transformers carry key/value states across segments as a persistent KV cache. This cache is often redundant across depth, with substantial inter-layer similarity in cached states Liu et al. [2024]. Motivated by this, inference-time studies on non-recurrent Transformers examine layer-wise KV-cache utility, showing it is heterogeneous across depth and making uniform budgeting suboptimal: SqueezeAttention assigns layer-wise KV budgets based on sensitivity Wang et al. [2025b], LAVa derives dynamic layer (and head) budgets from information-loss considerations Shen et al. [2025], and MiniCache exploits depth-wise redundancy by compressing highly similar KV states across adjacent middle-to-deep layers Liu et al. [2024].
Depth heterogeneity also appears in recurrent stacks, where long-range access is provided by stateful memories carried across segments. Notably, Rae et al. Rae and Razavi [2020] show that, for language modeling, better performance can be obtained by limiting the attention/memory range in lower layers, suggesting that long-range memory may be more valuable in higher layers. Our experiments on symbolic music yield the opposite pattern: allocating long horizons to lower layers outperforms allocating them to higher layers, indicating that the preferred depth placement is task-dependent.
Despite these findings, optimizing layer-wise horizon lengths—i.e., non-uniform per-layer memory windows that go beyond binary keep/drop decisions—remains largely unexplored in recurrent settings. DSMR addresses this gap by systematically varying per-layer horizons under a fixed total budget.
@@
2.4 Long-Context Symbolic Music Modeling under Resource Constraints
@@
A central goal in symbolic music generation is to capture long-range structure and full-piece coherence, since motifs, harmonic plans, and section-level returns can span thousands of events. To this end, music-specific long-context models explore structure-aware mechanisms for efficiency and long-term coherence, including improved relative attention Huang et al. [2019], fine–coarse attention over salient vs. nonsalient bars Yu et al. [2022], whole-song hierarchical generation using cascaded diffusion models Wang et al. [2024], and long-context pretraining for symbolic-music foundation models Qu et al. [2025]. LLM-style large-scale symbolic models further raise expectations on coherence and musicality Wang et al. [2025a].
These architectural advances expand the effective context available to the model. However, in practice, training pipelines in symbolic music generation remain dominated by efficiency-driven recipes: fixed-length segmentation, bounded-context truncation, or hierarchical/cascaded generation with fixed-scope local modeling, rather than end-to-end full-piece training. This parallels common LLM pretraining practice, which constructs fixed-length training sequences for efficiency; the specific sequence composition and masking strategy can meaningfully affect pretraining behavior Zhao et al. [2024], and Ding et al. Ding et al. [2024] show that “concatenate-then-split” can unnecessarily break documents that would otherwise fit within the context window, degrading coherent-content modeling. DSMR targets this gap by training on complete pieces end to end via streaming recurrence, avoiding the structural fragmentation inherent in fixed-length excerpt training.
3 Depth-Structured Music Recurrence (DSMR)
@@
Depth-Structured Music Recurrence (DSMR) is a training-time design for end-to-end full-piece symbolic music modeling via streaming recurrence. DSMR streams each composition left-to-right with stateful recurrent attention; under a fixed recurrent-state budget, it allocates depth-dependent (layer-wise) memory horizons to constrain the cost of carrying cross-segment states without reducing model depth. Our final instantiation is a two-scale DSMR schedule: long horizons in lower layers and a uniform short horizon elsewhere.
3.1 Recurrent backbone
@@
We adopt a Transformer-XL style segment-level recurrence. Each layer caches past key/value states and reuses them as additional attention context for subsequent segments. Unlike Transformer-XL, which discards older segments beyond a fixed window, DSMR ensures that at least one layer retains full-length (maximum-horizon) memory so that long-range context is never deterministically discarded; the specific layer(s) carrying long memory may differ across model variants. For completeness, details of Transformer-XL recurrence and its approximation to full self-attention are provided in Appendix A. Fig. 1 illustrates the segment-level recurrence with cached history and discarded context beyond the horizon.
@@
3.2 Training-time sampling and segmentation
@@
3.3 Full-attention approximation of a vanilla Transformer as a reference model
Transformer-XL enforces a fixed uniform horizon $m$ by retaining only the most recent $m$ cached states via the following memory update Dai et al. [2019], so attention to history is limited to a sliding window:
$M_{t}^{n}=\mathrm{SG}\!\left(\left[\,M_{t-1}^{n}\ ;\ H_{t-1}^{n}\,\right]\right)_{-m:},$ (1)
Appendix A.1 provides the exact stateful recurrent attention formulation for Transformer-XL, including memory concatenation, truncation, and stop-gradient across segments.
To approximate full-length context, DSMR extends Transformer-XL by setting the horizon to the maximum available prefix length for all layers, yielding a full-memory recurrent model.
Because DSMR uses truncated backpropagation through time (BPTT) with a stop-gradient across segments, it induces a biased approximation to full self-attention over the entire prefix, since the cached states are produced by parameters from earlier optimization steps. Let $\theta$ denote the current parameters, and let $\theta_{\text{old}}$ denote the parameters that produced the cached states. Then the recurrent attention uses history states computed by earlier parameters,
$(K_{\text{cache}},V_{\text{cache}})=\bigl(K(\theta_{\text{old}}),\ V(\theta_{\text{old}})\bigr),$ (2)